OpenStudy (anonymous):

Find the linear approximation if the independent variable changes from a to b. f(x) = x^4; a = 1, b = 0.98

OpenStudy (turingtest):

f(a)=1^4=1 f(b)=0.98^4=0.92 [f(b)-f(a)]/(b-a)=(0.92-1)/(0.98-1)=4

OpenStudy (anonymous):

The answer should be 0.92...

OpenStudy (anonymous):

I got 3.92 as the answer as well..